### v3

@@ 5 Threats to Validity
This section reports the potential threats to the validity of this research and its results, along with mitigation strategies that could help minimize the impacts of the outlined threats based on [17]. The threats are broadly classified across internal, construct, external, and conclusion validity.
@@ 5.1 Internal Validity
Internal validity examines the extent to which the study design, conduct, and analysis answer to the research questions without bias [17]. We discuss the following threats to internal validity.
@@
Instrument understandability. The participants of interviews and surveys may have a different understanding of the interview and survey instruments. Mitigation. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [29]. We also piloted both interview and survey instruments to ensure understandability (see Section 2.3.1 and Section 2.4.1). Our questionnaire for the survey were in English. However, during the interviews, we found that a few participants could not conveniently convey their answers in English. Therefore, we requested them to answer in their native languages, and for the latter we translated the answers into English.
@@ 5.2 Construct Validity
Construct validity focuses on whether the study constructs (e.g., interview protocol, survey questionnaire) are correctly defended [17]. Microservices issues, causes, and solutions are the core constructs of this study. Having said this, we identified the following threats.
@@ 5.3 External Validity
External validity refers to the extent to which the study findings could be broadly generalized in other contexts [17]. The sample size and sampling techniques might not provide a strong foundation to generalize the study results. It is the case for all three data collection methods (mining developer discussions from open-source microservices projects, interviews, and surveys) used in this study. Mitigation. (i) To minimize this threat, we derived the taxonomies of issues, causes, and solutions from a relatively large number of issues from 15 sampled open-source microservices projects belonging to different domains by involving multiple researchers. (ii) The taxonomies have been evaluated and improved by taking the feedback of experienced microservices practitioners through the interviews. (iii) A cross-sectional survey was conducted based on the derived and evaluated taxonomies. Overall, we received 150 valid responses from 42 countries of 6 continents (see Figure 3(a)) having varying experience (from less than one year to more than ten years, see Figure 3(b)) with different roles (see Figure 3(c)), and working with diverse domains (see Figure 3(d)) and programming languages and technologies (see Figure 3(e)) to develop microservices systems. Our study is primarily based on analyzing qualitative data, which presents potential threats, such as subjectivity bias, low reproducibility, data overload, and a lack of transparency that researchers may encounter. Mitigation. To mitigate these threats during data analysis, we employed thematic analysis by following the guidelines (see Section 2.2.2) and involved all the researchers to reduce subjectivity bias, enhance replicability, manage data overload, and ensure transparency.
@@
textitInterview data saturation: This threat arises from the possibility that the 15 interviews we conducted to validate and expand our proposed taxonomies may not be sufficient to achieve data saturation in our study. Mitigation. To address this threat, we conducted interviews with practitioners who had varied backgrounds (i.e., responsibilities, skills in programming languages, work domains, experience, countries) in microservices systems (see Table 2). Second, we asked open-ended questions, which allowed the interviewees to spontaneously express their views on the issues, causes, and solutions (see the Interview Questionnaire sheet in [19]). Third, we presented the interviewees with pre-existing taxonomies that were derived from analyzing developer discussions in 15 open-source microservices systems. We asked the interviewees to pinpoint any issue, cause, and solution they felt absent. Fourth, while we were analyzing interview data, we started observing recurring themes and patterns in the responses, which was an indication of data saturation. As we did not encounter any new significant issues, causes, or solutions after the 15th interview, we decided to stop conducting further interviews. However, we encourage further research involving more interviews or additional data sources to provide more insights and refinement to our findings.
@@ 5.4 Conclusion Validity
Conclusion validity is related to dealing with threats that affect the correct conclusions in empirical studies [17]. Concluded findings of the results may be based on a single author’s understanding and experience, and conflicts on the conclusions between authors may not be sufficiently discussed or resolved. Mitigation. To address this threat, the first author extracted and analyzed study data (i.e., from the developer discussions, interviews, and survey). All other authors comprehensively reviewed the data through multiple meetings. Conflicts on data analysis results were resolved through mutual discussions and brainstorming among all the authors. Different researchers can interpret the inclusion and exclusion criteria differently which influences the conclusions of the study. Mitigation. To minimize the effect of this issue, we applied the explicitly defined inclusion and exclusion criteria during the study data screening (see Section 2.2.2 - Step B). Finally, the interpreted results and conclusions have been confirmed by arranging several brainstorming sessions among all the authors.
@@ 6 Related Work
We performed a thorough literature search via eight major databases, including Google Scholar, IEEE Xplore, ACM Digital Library, ScienceDirect, Scopus, Web of Science, SpringerLink, and Wiley Online Library, to ensure that no relevant research study has been overlooked during the literature review. The literature search revealed that no study (except our previous work [18]) has empirically investigated the issues faced, causes reported, and solutions adopted by practitioners to develop microservices systems. We discuss the existing research that can be broadly classified into two categories including (i) mining OSS repositories to extract reusable knowledge for MSA (Section 6.1) and (ii) empirical studies on issues in microservices systems and software systems (Section 6.2). A conclusive summary and comparative analysis (Section 6.3) position the proposed research in the context of mining issues from OSS repositories and justify the scope and contributions of this research.
@@ 6.1.1 Mining (Anti-)Patterns and Tactics to Architect MSA
From an MSA perspective, architectural patterns [34] and tactics [66] represent empirically derived knowledge, best practices, and recurring solutions to address frequently occurring issues during architecture-centric design and development of service-driven systems [10]. To discover patterns and tactics for architecting MSA, a number of studies have focused on analyzing open source repositories – mining version controls, searching change logs, and exploring design documents etc. – to investigate historical data that represents recurring solutions as patterns [10, 67, 68, 69].
@@
The investigation of historical data involves postmortem analysis of development activities (e.g., software refactoring, testing, evolution) [70] as well as the knowledge of architects (e.g., developer discussions, code documentation) [71], which is captured via open source repositories, such as GitHub, Stack Overflow, or customized databases [10, 67, 71]. Specifically, Marquez et al. [10] explored source code artifacts of microservices projects, i.e., configuration files, and framework dependencies, available on GitHub to mine 17 architectural patterns addressing a set of quality attributes. Compared to the design and implementation-specific knowledge for microservices reported in [10, 67], Armin et al. [68] explored the evolution phase from traditional architectures to MSA and identified 15 migration patterns that support the evolution of legacy software towards a modular MSA. Compared to the pattern-based solutions discussed above, architectural tactics represent design decisions that focus on improving one specific quality aspect of MSA, such as service availability and fault avoidance for security critical systems [66]. Amoroso et al. [72] curated and manually labeled a dataset of 378 Dockerized open-source microservice projects from an initial pool of 389,559 candidates. The dataset includes project metadata such as contributors, dependencies, and application purposes, supporting future empirical research on microservice practices and evolution. Lercher et al. [73] investigated API evolution in microservice systems through interviews with practitioners, identifying six strategies and six challenges in handling REST API changes. They proposed automated mechanisms for extracting structural and behavioral API changes and notifying consumer teams. Their findings highlight the socio-technical complexities of API evolution in microservices, complementing prior work on migration and evolvability. Lubas et al. [74] mined GitHub to curate two datasets: one of 553 microservice applications and another of 8 representative workload repositories. Their analysis provides insights into programming languages, frameworks, and architectural components used in practice, supporting benchmarking and empirical studies on workload characterization in microservice systems.
While patterns promote best practices to develop MSA, anti-patterns represent a class of patterns that have been perceived as best practices and commonly used but are proven to be ineffective and/or counterproductive, such as bad smells in service design [14]. Taibi et al. [70] conducted interviews of microservices developers, i.e., accumulating practitioners’ experiences to create a taxonomy of 20 anti-patterns including organizational (team oriented and technology/tool oriented) and technical (internal and communication) anti-patterns. The anti-pattern taxonomy aims to help microservices practitioners to avoid counterproductive patterns and design decisions. Tighilt et al. [75] proposed a specification and detection approach for microservice antipatterns. Their study defined formal detection rules, validated them on open-source systems, and demonstrated their utility for supporting maintenance and quality improvement in microservice architectures. This work goes beyond descriptive taxonomies by providing automated mechanisms for identifying recurring bad practices. In comparison to the repository mining approaches that investigate code and architectural-centric artifacts available on GitHub such as [66, 10, 69], Bandeira et al. [71] reflected a human-centric view of MSA design by analyzing developer discussions on Stack Overflow. Their study classified a total of 1,043 microservice tagged posts into three categories namely technical (44.87%), conceptual (30%), and non-related (25.13%) discussions to explore the processes, tools, issues, solutions, etc. that developers find most exciting or challenging while implementing MSA.
@@ 6.1.2 Analyzing Issues for Evolving Systems to MSA
In recent years, research and practices on the migration of legacy systems to microservices systems have gained significant attention with empirical evidence derived from industrial practices on the role of MSA in software evolvability [76, 40], service extraction from legacies [77, 78], and identifying the motivations and challenges for legacy migration to MSA [2]. For instance, Bogner et al. [76] reported the results of 17 structured interviews with 14 microservices practitioners from 10 projects to investigate a multitude of issues, such as tool support, patterns, and process to manage TD and enhance software evolvability using microservices. The findings of the interviews recommended a number of techniques, including but not limited to code review and service slicing, that can address a number of microservices issues related to service granularity, composition, coupling, and cohesion. Lenarduzzi et al. [40] explored the issues of TD when legacy software is migrated to microservices. They investigated four years history of a project having 280K lines of code and concluded that although TD spiked initially due to the development of new microservices, however; after a short period of time the TD grew slower in the microservices system than in the (legacy) monolithic system. Carvalho et al. [78] conducted an online survey with 26 specialists, followed by individual interviews with 7 of them to understand the challenges pertaining to the migration of existing systems to microservices architecture. Their study results highlight that extracting the microservices from legacy components and monolithic source code modules represents the most critical challenge during the re-engineering or migration of legacies toward microservices systems. A number of empirical studies, such as [77, 2], engaged industrial practitioners to understand the processes, motivations, and challenges related to legacy system migration in general and service extraction in particular. The results of these studies provide recommendations and guidelines about service maintainability and scalability as the most important motivations for the enterprise-scale adoption of MSA.
@@ 6.2 Empirical Studies on Issues in Microservices Systems and Software Systems
6.2.1 Smells & Performance Issues in MSA Systems
Code smells and architectural smells (a.k.a bad smells) are often synonyms as a microservices anti-pattern reflecting symptoms of poorly designed microservices that decrease code understandability and increase effort for maintainability [70, 14]. According to practitioners’ suggestions and industrial case studies, detecting bad smells in microservices systems is critical for large-scale microservice systems [14]. Walker et al. [79] provided tool support for the automatic detection of bad smells in microservices systems, and the tool MSANose can detect up to eleven microservices specific bad smells within microservices applications using bytecode and/or source code analysis throughout the development process or even before its deployment to production. In addition, a number of other prevalent issues in microservices systems, such as faults, bugs, performance issues, and service decomposition problems, are detailed in [15, 16, 80]. More specifically, Wu et al. [16] proposed a solution named Microservice Root Cause Analysis (MicroRCA), which works by inferring the root causes of performance issues by correlating application performance symptoms with corresponding system resource utilization. Their proposed solution MicroRCA can addresses performance related issues in microservices systems by analyzing resource utilization and throughput of the services.
@@ 6.2.2 Taxonomies of Issues & Faults in Software Systems
While exploring issues from the microservices system point of view, it is important not to overlook the most recent taxonomies, empirical studies, and proposed solutions that address a multitude of issues, errors, and faults in non-MSA systems such as, deep learning [81] and application build systems [45]. In particular, a taxonomy of the types of faults in deep learning systems [81] and the types of build issues, their symptoms, and fix patterns [45] inspired our work on microservices issues, causes, and solutions. From the MSA perspective, a recently conducted study [15] identified typical faults occurring in microservices systems, practices of service debugging, and challenges faced by developers while addressing these faults. For example, a fault such as “transactional service failure” is due to overloaded requests to a third-party (payment gateway) service, ultimately leading to denial of service issues. Our proposed research draws inspiration from empirically derived taxonomies [81, 45] and goes beyond issue categorization to investigate their causes and proposed solutions as resolution strategies to fix multi-faceted issues related to Security, Testing, and Configuration that impact architecting and implementing microservices systems.

### v4

@@ 5 Threats to Validity
This section reports the potential threats to the validity of this research and its results, along with mitigation strategies that could help minimize the impacts of the outlined threats based on [easterbrook2008selecting]. The threats are broadly classified across internal, construct, external, and conclusion validity.
@@ 5.1 Internal Validity
Internal validity examines the extent to which the study design, conduct, and analysis answer to the research questions without bias [easterbrook2008selecting]. We discuss the following threats to internal validity.
@@
Instrument understandability. The participants of interviews and surveys may have a different understanding of the interview and survey instruments. Mitigation. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [kitchenham2008personal]. We also piloted both interview and survey instruments to ensure understandability (see Section 2.3.1 and Section 2.4.1). Our questionnaire for the survey were in English. However, during the interviews, we found that a few participants could not conveniently convey their answers in English. Therefore, we requested them to answer in their native languages, and for the latter we translated the answers into English.
@@ 5.2 Construct Validity
Construct validity focuses on whether the study constructs (e.g., interview protocol, survey questionnaire) are correctly defended [easterbrook2008selecting]. Microservices issues, causes, and solutions are the core constructs of this study. Having said this, we identified the following threats.
@@ 5.3 External Validity
External validity refers to the extent to which the study findings could be broadly generalized in other contexts [easterbrook2008selecting]. The sample size and sampling techniques might not provide a strong foundation to generalize the study results. It is the case for all three data collection methods (mining developer discussions from open-source microservices projects, interviews, and surveys) used in this study. Mitigation. (i) To minimize this threat, we derived the taxonomies of issues, causes, and solutions from a relatively large number of issues from 15 sampled open-source microservices projects belonging to different domains by involving multiple researchers. (ii) The taxonomies have been evaluated and improved by taking the feedback of experienced microservices practitioners through the interviews. (iii) A cross-sectional survey was conducted based on the derived and evaluated taxonomies. Overall, we received 150 valid responses from 42 countries of 6 continents (see Figure 3(a)) having varying experience (from less than one year to more than ten years, see Figure 3(b)) with different roles (see Figure 3(c)), and working with diverse domains (see Figure 3(d)) and programming languages and technologies (see Figure 3(e)) to develop microservices systems. Our study is primarily based on analyzing qualitative data, which presents potential threats, such as subjectivity bias, low reproducibility, data overload, and a lack of transparency that researchers may encounter. Mitigation. To mitigate these threats during data analysis, we employed thematic analysis by following the guidelines (see Section 2.2.2) and involved all the researchers to reduce subjectivity bias, enhance replicability, manage data overload, and ensure transparency.
@@
textitInterview data saturation: This threat arises from the possibility that the 15 interviews we conducted to validate and expand our proposed taxonomies may not be sufficient to achieve data saturation in our study. Mitigation. To address this threat, we conducted interviews with practitioners who had varied backgrounds (i.e., responsibilities, skills in programming languages, work domains, experience, countries) in microservices systems (see Table 2). Second, we asked open-ended questions, which allowed the interviewees to spontaneously express their views on the issues, causes, and solutions (see the Interview Questionnaire sheet in [replpack]). Third, we presented the interviewees with pre-existing taxonomies that were derived from analyzing developer discussions in 15 open-source microservices systems. We asked the interviewees to pinpoint any issue, cause, and solution they felt absent. Fourth, while we were analyzing interview data, we started observing recurring themes and patterns in the responses, which was an indication of data saturation. As we did not encounter any new significant issues, causes, or solutions after the 15th interview, we decided to stop conducting further interviews. However, we encourage further research involving more interviews or additional data sources to provide more insights and refinement to our findings.
@@ 5.4 Conclusion Validity
Conclusion validity is related to dealing with threats that affect the correct conclusions in empirical studies [easterbrook2008selecting]. Concluded findings of the results may be based on a single author’s understanding and experience, and conflicts on the conclusions between authors may not be sufficiently discussed or resolved. Mitigation. To address this threat, the first author extracted and analyzed study data (i.e., from the developer discussions, interviews, and survey). All other authors comprehensively reviewed the data through multiple meetings. Conflicts on data analysis results were resolved through mutual discussions and brainstorming among all the authors. Different researchers can interpret the inclusion and exclusion criteria differently which influences the conclusions of the study. Mitigation. To minimize the effect of this issue, we applied the explicitly defined inclusion and exclusion criteria during the study data screening (see Section 2.2.2 - Step B). Finally, the interpreted results and conclusions have been confirmed by arranging several brainstorming sessions among all the authors.
@@ 6 Related Work
We performed a thorough literature search via eight major databases, including Google Scholar, IEEE Xplore, ACM Digital Library, ScienceDirect, Scopus, Web of Science, SpringerLink, and Wiley Online Library, to ensure that no relevant research study has been overlooked during the literature review. The literature search revealed that no study (except our previous work [waseem2021nature]) has empirically investigated the issues faced, causes reported, and solutions adopted by practitioners to develop microservices systems. We discuss the existing research that can be broadly classified into two categories including (i) mining OSS repositories to extract reusable knowledge for MSA (Section 6.1) and (ii) empirical studies on issues in microservices systems and software systems (Section 6.2). A conclusive summary and comparative analysis (Section 6.3) position the proposed research in the context of mining issues from OSS repositories and justify the scope and contributions of this research.
@@ 6.1.1 Mining (Anti-)Patterns and Tactics to Architect MSA
From the MSA perspective, patterns [R1] and tactics [A66-genfer2025understanding] represent empirically grounded knowledge, that promote reuse of best practices, as recurring solutions to address frequently occurring issues during architecture-centric development of service-driven software systems [A10-daniel2023towards]. To derive empirically-grounded reuse knowledge, i.e., discovering patterns and tactics for architecting MSA, a number of studies such as [A10-daniel2023towards, A67-soldani2021mutosca, A68-fritzsch2022towards, A69-bacchiega2022microservices] have focused on analyzing open-source code and project repositories such as mining version controls, searching change logs, and exploring design documents etc. to investigate historical data that can reveal recurring solutions as patterns. The investigation of historical data involves postmortem analysis to detect anti-patterns and bad smells [A70-cerny2023catalog] and mining repositories such as GitHub [A71-beahan2025characterizing] to identify vulnerabilities in MSA systems. Daniel et al. [A10-daniel2023towards] employed a metric-based approach and used two case studies to discover microservice patterns, while also evaluating the accuracy and applicability of the approach based on developers’ feedback. In a similar work, Soldani et al. [A67-soldani2021mutosca] have developed a TOSCA toolchain to detect and resolve architectural smells and to ensure that complex microservice ecosystems remain aligned with core design principles. Compared to the metric and tool-based approaches for mining MSA specific knowledge, Fritzsch et al. [A68-fritzsch2022towards] employed a mixed-method approach, synthesising the findings of literature review and industry expert interviews to derive a framework for architectural refactoring and migration in the context of microservice systems.
@@
Compared to the pattern-based solutions discussed above, architectural tactics represent design decisions that focus on improving one specific quality aspect of MSA, such as service availability and fault avoidance for security critical systems [A66-genfer2025understanding] . Amoroso et al.[A72-amoroso2024dataset] curated and manually labeled a dataset of 378 Dockerized open-source microservice projects from an initial pool of 389,559 candidates. The dataset includes project metadata such as contributors, dependencies, and application purposes, supporting future empirical research on microservice practices and evolution. Lercher et al. [lercher2024managing] investigated API evolution in microservice systems through interviews with practitioners, identifying six strategies and six challenges in handling REST API changes. They proposed automated mechanisms for extracting structural and behavioral API changes and notifying consumer teams. Their findings highlight the sociotechnical complexities of API evolution in microservices, complementing prior work on migration and evolvability. Lubas et al. [lubas2025microservice] mined GitHub to curate two datasets: one of 553 microservice applications and another of 8 representative workload repositories. Their analysis provides insights into programming languages, frameworks, and architectural components used in practice, supporting benchmarking and empirical studies on workload characterization in microservice systems.
While patterns promote best practices to develop MSA, anti-patterns represent a class of patterns that have been perceived as best practices and commonly used but are proven to be ineffective and/or counterproductive, such as bad smells in service design [A14-avritzer2025architecture]. Cerny et al. [A70-cerny2023catalog] conducted a tertiary study synthesising secondary studies on poor design practices in microservice architectures – referred to as anti-patterns. The study systematically identifies recurring anti-patterns and bad smells, classifies them into coherent categories, and summarizes existing approaches for their detection. The results offer a structured reference for developers to design higher-quality microservice systems and for researchers seeking to assess system quality through anti-patterns. Tighilt et al. [tighilt2023maintenance] proposed a specification and detection approach for microservice anti-patterns. Their study defined formal detection rules, validated them on open-source systems, and demonstrated their utility for supporting maintenance and quality improvement. This work goes beyond descriptive taxonomies by providing automated mechanisms for identifying recurring bad practices. In the context of repository-based knowledge mining approaches that investigate code and architectural-centric artifacts available on GitHub such as [A2-michael2023empirical, A10-daniel2023towards, A66-genfer2025understanding, A69-bacchiega2022microservices], Beahan et al. [A71-beahan2025characterizing] empirically analyzed security vulnerabilities in microservice-based systems by examining 30 open-source projects using multiple detection tools. The results indicate that vulnerabilities primarily arise from application code, dependencies, and container configurations. While microservices do not introduce fundamentally new vulnerability types, the findings of the study highlight the need for stronger practices in dependency management, container configuration, and secure data handling.
@@ 6.1.2 Analyzing Issues for Evolving Systems to MSA
In recent years, research and practices on the evolution of microservices systems have gained significant attention based on analysis of methods and techniques that enable a systematic evolution of MSAs [A40-cerny2025analyzing, A76-nogueira2024insights], service extraction from legacies [A77-mohottige2025reengineering, A78-abdellatif2025identifying], and identifying the motivations and challenges for legacy migration to MSA [A2-michael2023empirical]. For example, Nogueira et al [A76-nogueira2024insights] investigated a practitioner-led migration from monolithic legacy systems to microservice architectures, focusing on motivations, activities, and data consistency strategies. The study conducted a questionnaire-based survey with 53 participants and highlighted that while technical benefits like scalability and maintenance drive adoption, significant challenges remain in testing, monitoring, and managing decentralized data consistency. Beyond the practitioners’ survey-based studies to enable and enhance legacy migration to service-driven systems, Abdellatif et al [A78-abdellatif2025identifying] evaluated the transition of legacy systems to service-oriented architectures by comparing academic and industrial service identification approaches. The comparison identified gaps between the two. To bridge this gap, the authors introduce ServiceMiner, a bottom-up identification approach that utilizes source-code analysis and structural patterns to detect reusable functionalities when documentation is unavailable. Several recently published empirical studies, such as [A2-michael2023empirical, A66-genfer2025understanding, A76-nogueira2024insights] have focused on experimental analysis, case studies, and practitioners’ feedback to understand the processes, motivations, and challenges related to legacy system migration in general and service extraction in particular. The results of these studies provide empirically-grounded recommendations and guidelines related to service design, maintainability and scalability as the prime motivations for the enterprise-scale adoption of MSAs.
@@ 6.2 Empirical Studies on Issues in Microservices Systems and Software Systems
6.2.1 Detecting Bad Smells and Performance Issues
Code smells and architectural smells, often collectively referred to as bad smells, represent an anti-pattern that reflect symptoms of poorly designed microservice system that decrease code understandability and increase effort for maintainability [70, 14]. Bad smells represent low-level design problems and poor coding decisions that often indicate the presence of anti-patterns with a negative impact on software quality [A14-avritzer2025architecture]. In absence of a systematic process and tool support for automation, smell detection can be a laborious task that is time-consuming and error-prone. In the context of automating the detection of bad smells, Monsef et al. [A79-monsef2025detecting] have proposed an automated detection approach using graph neural networks, modeling microservices as graphs to identify four architectural anti-patterns. To mitigate data scarcity, large language models are employed to generate and augment architectural graph datasets. In addition, a number of other prevalent issues in microservices systems, such as faults, bugs, performance issues, and service decomposition problems, are detailed in [A15-zhang2024trace, A80-wizenty2023model]. More specifically, Zhang et al.[A15-zhang2024trace] have proposed an approach namely TraceContrast that applies a trace-based root cause localization to address the complexity of microservice failures by exploiting fine-grained distributed traces. Empirical evaluation on a standard microservice benchmark demonstrates that TraceContrast achieves significantly higher accuracy than existing methods at both multidimensional and instance-level localization of microservice faults. The work by Wizenty et al. [A80-wizenty2023model] tackles the challenge of automated detection of security bad smells and apply refactoring of microservice systems using a model-driven detection approach.
@@ 6.2.2 Taxonomies of Issues & Faults in Software Systems
While exploring issues from the microservices system point of view, it is important not to overlook the most recent taxonomies, empirical studies, and proposed solutions that address a multitude of issues, errors, and faults in non-MSA systems such as, deep learning-based bug classification [Humbatova20] and application build systems [A45-lou2020understanding]. In particular, a deep learning approach to classify bugs for code inspection [Humbatova20] and the types of build issues, their symptoms, and fix patterns [A45-lou2020understanding] also inspired our work on microservices issues, causes, and solutions. From the MSA perspective, a recently conducted study [A15-zhang2024trace] focuses on localization of root cause analysis to resolve the issues related to microservice bad smells. Our proposed research draws inspiration from empirically derived taxonomies [Humbatova20, A45-lou2020understanding] and goes beyond issue categorization to investigate their causes and proposed solutions as resolution strategies to fix multi-faceted issues related to Security, Testing, and Configuration that impact architecting and implementing microservices systems.